# thanks: Corresponding author

### v1

@@ II.4 Natural Auxiliary Functions (NAF)
Analogous to the reduction of the orbital basis set using Frozen Natural Orbitals (FNO), the dimensionality of the auxiliary basis can be effectively reduced through the use of Natural Auxiliary Functions (NAFs). [kallaySystematicWayCost2014] The NAFs are generated using a singular value decomposition of the three-centered two-electron matrix $\mathbf{J}$ in the molecular orbital basis.
@@
Where $\bm{\Sigma}$ denotes the diagonal matrix containing singular values and $\mathbf{A}$ and $\mathbf{B}$ denote the left and right singular vectors, respectively. Alternatively, one can obtain them by diagonalizing the matrix
@@
Following the recommendation of Kallay and co-workers [mesterReducedcostLinearresponseCC22017], the three-center two-electron integral of virtual virtual type ($\mathbf{J}_{ab}$) has been used to construct $\mathbf{W}$. One can discard NAFs corresponding to the eigenvalues lower than a particular threshold. The rest of the eigenvectors are collected in the matrix $\tilde{\mathbf{N}}$. The matrix $\mathbf{J}$ can be truncated by transforming the auxiliary function index of the matrix $\mathbf{J}$ as
@@
The $\tilde{\mathbf{J}}$ are subsequently used to generate the integrals.
@@ II.5 Correction for the truncation
Second-order perturbative energy correction has been shown to improve the accuracy of truncated natural orbital based wave-function methods. [kallaySystematicWayCost2014, mukhopadhyayStatespecificFrozenNatural2023, mannaEfficientStateSpecificNatural2024] In this study, we have included a correction for the natural orbital based truncation to the SS-FNO-EA-ADC(3) results.
@@
Here, the difference between EA-ADC(2) results in the canonical and SS-FNO basis is added to the uncorrected third-order EA-ADC results as a correction for the natural orbital and natural auxiliary function truncations.
@@ III Computational Details
The SS-FNO-EA-ADC(3) method is implemented in BAGH, [duttaBAGHQuantumChemistry2025], our in-house quantum chemistry software package. BAGH is primarily written in Python, with computationally expensive parts optimized using Cython and Fortran. BAGH is currently integrated with four external packages for the generation of integrals, namely, PySCF [sunLibcintEfficientGeneral2015, sunPySCFPythonbasedSimulations2018, sunRecentDevelopmentsPySCF2020], GAMESS-US [barcaRecentDevelopmentsGeneral2020], DIRAC [bastDIRACProgramAtomic2023] and socutils [wangXubwaSocutils2025]. The implementation of the SS-FNO-EA-ADC(3) method presented in this manuscript utilizes the PySCF interface for integral generation. The steps involved in the SS-FNO-EA-ADC(3) method are as the followings:
@@
    After a successful SCF convergence, the three-centered two-electron integrals $(P|ij)$, $(P|ab)$ and $(P|ia)$ are formed.
@@
    MP2 calculation is performed in the canonical basis, and the one-particle reduced density $\left(\mathbf{D}_{ab}^{MP2}\right)$ is formed in the canonical basis.
@@
    EA-ADC(2) calculation is performed in the canonical basis.
@@
    Looping over each root:
@@
      The virtual-virtual block of the EA-ADC(2) reduced density $\left(\mathbf{D}_{ab}^{EA-ADC(2)}\right)$ for the respective root is formed.
@@
      EA-ADC(2) reduced density is added to MP2 one-particle reduced density to form state-specific reduced density.
@@
      The state-specific reduced density is diagonalized, and the natural orbitals are sorted based on occupancies.
@@
      The virtual natural orbitals having occupancies lower than the SS-FNO threshold are dropped off for further calculations.
@@
      EA-ADC(2) and EA-ADC(3) calculations are performed in the truncated natural orbital basis and the corresponding truncated NAF basis.
@@
      A perturbative correction taking the difference of the canonical and the truncated EA-ADC(2) results is added to the EA-ADC(3) result.
A schematic diagram on the algorithm of the SS-FNO-EA-ADC(3) method is presented in FIG. 1
All calculations in this work are performed using the density-fitting approximation. Core electrons are kept frozen for all calculations.
For the Complete Basis Set (CBS) extrapolation of the electron affinity value, the two-point cubic extrapolation scheme, introduced by Helgaker and coworkers, has been employed. [helgakerBasissetConvergenceCorrelated1997, halkierBasissetConvergenceCorrelated1998]
@@
where $\alpha$ is the parameter and $E_{CE}^{CBS}$ is the correlation energy at the CBS limit. Here, CE stands for correlation energy. We have extrapolated our aug-cc-pVXZ (X=D,T) electron affinity results to the CBS limit using Eq. (33). The el
@@ IV.1 Optimization of thresholds
To assess the performance of the SS-FNO scheme for the electron-attachment problem, we examined the deviation in EA-ADC(3) electron affinity values as a function of the virtual space size, both in the canonical molecular orbital (MO) basis and in the natural orbital bases (FNO and SS-FNO) for Ozone (see FIG. 2(a)). The experimental geometry of the Ozone molecule [johnsonComputationalChemistryComparison2002] is used for the calculations. The aug-cc-pVQZ basis set and the aug-cc-pVQZ/C basis set have been used for the calculations. No NAF-based truncation has been used for the calculations. The electron affinities (eV) at different percentages of virtual orbitals are provided in the Supplementary Material (see TABLE S1).
@@
The truncation based on occupation number is more robust than the percentage of virtual orbitals retained in the natural orbital-based implementation of the wave-function method [kallaySystematicWayCost2014]. In occupation number-based truncation schemes, the virtual natural orbitals up to a certain truncation threshold have been included in the calculations. FIG. 2(b) presents the convergence of error at different natural orbital-based truncations of the EA-ADC(3) method with respect to the untruncated canonical EA-ADC(3) method. It can be seen that the SS-FNO-based scheme converges more rapidly than the corresponding FNO-based one. The EA values in the SS-FNO-based scheme converge at the threshold $10^{-5}$, while the FNO-based scheme shows a large error of -0.641 eV at the same threshold. The FNO-based scheme eventually converges at the threshold of $10^{-7}$. It shows that the SS-FNO framework is a more suitable tool for reduced-cost EA-ADC(3) calculations. The inclusion of the perturbative correction significantly enhances the accuracy, and the error converges at the threshold of $10^{-4}$. At this threshold, SS-FNO-EA-ADC(3) selects only $38.6\%$ of the total virtual space. Further reduction in computational cost can be achieved by reducing the auxiliary basis dimension using Natural Auxiliary Functions (NAFs). Fig. 3 illustrates the convergence of the error (in eV) in the EA value of O3 relative to the canonical reference, with decreasing NAF threshold and a fixed SS-FNO threshold of $10^{-4}$ in the EA-ADC(3) method. The EA values (in eV) at different NAF thresholds and the reduced dimension of the auxiliary basis are tabulated in TABLE S3. The aug-cc-pVQZ basis set and the aug-cc-pVQZ/C auxiliary were used for the calculations. The convergence plot suggests that even at the NAF threshold of $10^{-1}$, the EA value converges with an error of 0.022 eV with only around $70\%$ of the auxiliary basis functions retained. With the inclusion of perturbative corrections, the EA values show a negligible error (<0.01<0.01 eV) at the threshold of $10^{-1}$. However, for further benchmarking of different closed-shell systems, a conservative NAF threshold of $10^{-2}$ has been chosen.
@@ IV.2 Benchmarking on EA24 test set
To further assess the performance of the SS-FNO ADC(3) method, we have calculated the electron affinity value of the molecule included in the EA24 test set of Scherrill and co-workers [richardAccurateIonizationPotentials2016]. It consists of a group of 24 organic acceptor molecules, which contain bound electron-attached states of valence type, and experimental values are available for most of them. The electron affinity values of the molecules in EA24 test set have been extensively studied and benchmarked for EA due to its practical implications in organic photovoltaics. [richardAccurateIonizationPotentials2016, gallandiAccurateIonizationPotentials2016, knightAccurateIonizationPotentials2016, dolgounitchevaCorrectionAccurateIonization2017, duttaDomainbasedLocalPair2019, shaalanalagAccuratePredictionVertical2022, behjouElectronAffinitiesEquationofMotion2025, opokuInitioElectronPropagators2024, opokuNewGenerationElectronPropagatorMethods2024, mesterVerticalIonizationPotentials2023, galynskaExploringElectronAffinities2024] The molecules involved in the EA24 test set are shown in FIG. 4, and the corresponding geometries of the molecules are taken from Ref. knightAccurateIonizationPotentials2016. A few of the molecules have more than one bound state. However, we have considered only the lowest energy electron-attached state for the sake of simplicity.
@@
We have calculated the vertical electron attachment (EA) energies of these molecules at three truncation thresholds, $10^{-3}$, $10^{-4}$, and $10^{-5}$, using the aug-cc-pVDZ basis set in conjunction with the aug-cc-pVDZ/C auxiliary basis set. Figure 5 presents the distribution of errors with respect to the canonical EA-ADC(3) results. The plot shows a smooth convergence of the EA values as the truncation threshold is tightened. Furthermore, the inclusion of perturbative corrections leads to a noticeable reduction in both the spread and magnitude of the errors. The corresponding statistical parameters are summarized in Table 1, where the following measures are employed: MAD (Maximum Absolute Deviation), ME (Mean Error), MAE (Mean Absolute Error), STD (Standard Deviation), and RMSD (Root Mean Squared Deviation).The MAD value decreases from 0.378 eV at the $10^{-3}$ threshold to 0.042 eV at $10^{-4}$, and further to 0.003 eV at $10^{-5}$. The inclusion of the perturbative correction substantially improves the results, particularly at higher thresholds. For instance, at the $10^{-3}$ threshold, the MAD reduces from 0.378 eV to 0.170 eV upon applying the perturbative correction, while at $10^{-4}$, it decreases from 0.042 eV to 0.025 eV. At the tightest threshold ($10^{-5}$), the perturbative correction has a negligible effect, indicating convergence with respect to the SS-FNO truncation. The ME values suggest that the SS-FNO truncation systematically underestimates the EA, as reflected by the identical magnitudes of ME and MAE. All statistical parameters become negligible at the $10^{-5}$ threshold, confirming near-complete convergence of the SS-FNO approach.
Table 1: Statistical analysis of errors (in eV) of $S$-FNO-EA-ADC(3) with respect to the canonical results at different FNO thresholds.
@@
Figure 6: Percentage of retained VNOs and absolute errors in EA values (eV) relative to canonical results, as a function of system size, for EA24 molecules using the aug-cc-pVDZ basis and aug-cc-pVDZ/C auxiliary basis
FIG. 6 presents the percentage of virtual natural orbitals (VNOs) retained (blue) and the absolute error in the EA energy (pink) as a function of the system size. The percentage of retained SS-FNOs exhibits a variation ranging from approximately 65% to 85% across different systems. The truncation introduces a nearly uniform error across the test set, with the maximum absolute deviation remaining below 0.025 eV (including the perturbative correction). It demonstrates that the SS-FNO method can select an appropriate active space regardless of system size. The truncations are even larger at large basis sets. For example, the average truncation for the EA24 test set is $\approx$ 52 $\%$ in the aug-cc-pVTZ basis set (See Table S7). We have also performed the statistical analysis of error for the NAF truncation. FIG. 7 provides a statistical error distribution compared to the canonical result for a systematic comparison of the SS-FNO-EA-ADC(3) framework evaluated at various NAF truncations for two NAF thresholds ($10^{-1}$ and $10^{-2}$). The perturbative correction is included for this comparison.
@@
Table 3: The electron affinities (in eV) of the first electron attached states of EA24 test set molecules calculated using various EA-ADC methods in CBS limit (in the SS-FNO basis, including perturbative correction) and compared with CCSD(T) results. The EA-EOM-DLPNO-CCSD/CBS results are also shown for comparison. The experimental values are provided for reference. The SS-FNO threshold is kept at $10^{-4}$ and NAF threshold at $0^{-2}. 11$footnotetext: NormalPNO with TCutPNOSingles 1e-9 is used.
Molecule ADC(2)111in the canonical basis. $S$-FNO-ADC(3) $S$-FNO-sm-ADC[(2)+x(3)] EOM-DLPNO-CCSD22footnotemark: 2 CCSD(T) [richardAccurateIonizationPotentials2016] Expt. [informaticsNISTChemistryWebBook]
@@
TABLE 3 summarizes the electron affinities (EAs) of the lowest electron-attached states for all molecules in the EA24 test set, computed using various EA-ADC methods and compared with the CCSD(T) reference values extrapolated to the complete basis set (CBS) limit. [richardAccurateIonizationPotentials2016] The CBS extrapolation was performed using results obtained with the aug-cc-pVXZ (X = D, T) basis sets. For the second-order ADC calculations, canonical results are reported, whereas for the third-order ADC, only the corrected SS-FNO values are considered for comparison. The ADC(2) method exhibits the largest deviation from the CCSD(T) benchmarks, with a mean absolute deviation (MAD) of 0.973 eV. Inclusion of third-order corrections substantially improves the accuracy, as reflected in the reduced MAD of 0.56 eV for the SS-FNO-ADC(3) method, although this value remains larger than that observed for the corresponding EOM-DLPNO-CCSD method. The SS-FNO-sm-ADC[(2)+x(3)] approach demonstrates significantly improved performance, yielding a MAD of only 0.26 eV and showing close agreement with the EOM-DLPNO-CCSD results, which exhibit slightly smaller MAD and MAE but a larger standard deviation (STD). Although a direct comparison with experimental data is less meaningful due to the omission of vibrational contributions in the calculations, the SS-FNO-sm-ADC[(2)+x(3)] method at the CBS limit nonetheless exhibits marginally better agreement with experiment than the corresponding CCSD(T) results.

### v2

@@ II.4 Natural Auxiliary Functions (NAF)
Analogous to the reduction of the orbital basis using frozen natural orbitals (FNOs), the dimensionality of the auxiliary basis can be reduced using natural auxiliary functions (NAFs). [30] NAFs are generated using a singular value decomposition of the three-center two-electron matrix $\mathbf{J}$ in the molecular orbital basis.
@@
Here $\bm{\Sigma}$ denotes the diagonal matrix containing singular values and $\mathbf{A}$ and $\mathbf{B}$ denote the left and right singular vectors, respectively. Alternatively, the NAFs can be obtained by diagonalizing the matrix
@@
Following the recommendation of Kállay and co-workers [42], the virtual–virtual block of the three-centered integral matrix $\mathbf{J}_{ab}$ is used to construct $\mathbf{W}$. NAFs corresponding to eigenvalues smaller than a chosen threshold are discarded. The remaining eigenvectors are collected in the matrix $\tilde{\mathbf{N}}$. The matrix $\mathbf{J}$ is truncated by transforming its auxiliary index as
@@
The $\tilde{\mathbf{J}}$ are subsequently used in the generation of molecular integrals.
@@ II.5 Correction for the truncation
Second-order perturbative energy correction has been shown to improve the accuracy of truncated natural orbital-based wave-function methods. [30, 46, 38] In this study, we have included a correction for the natural orbital-based truncation to the SS-FNO-EA-ADC(3) results.
@@
Here, the difference between the EA-ADC(2) results in the canonical and SS-FNO bases is added to the uncorrected EA-ADC(3) result as a correction for both natural orbital and auxiliary basis truncations.
@@ III Computational Details
The SS-FNO-EA-ADC(3) method is implemented in BAGH, [16], our in-house quantum chemistry software package. BAGH is primarily written in Python, with computationally intensive components optimized using Cython and Fortran. BAGH is currently integrated with four external packages for the generation of integrals, namely, PySCF [71, 69, 70], GAMESS-US [3], DIRAC [5] and socutils [79]. The implementation of SS-FNO-EA-ADC(3) described in this work uses the PySCF interface for integral generation. The steps involved in the SS-FNO-EA-ADC(3) method are as follows:
@@
    After a successful SCF convergence, the three-centered two-electron integrals $(P|ij)$, $(P|ab)$ and $(P|ia)$ are generated.
@@
    An MP2 calculation is performed in the canonical basis, and the one-particle reduced density $\left(\mathbf{D}_{ab}^{MP2}\right)$ is constructed.
@@
    An EA-ADC(2) calculation is performed in the canonical basis.
@@
    The following steps are performed for each root:
@@
      The virtual-virtual block of the EA-ADC(2) reduced density $\left(\mathbf{D}_{ab}^{EA-ADC(2)}\right)$ is constructed for the corresponding root.
@@
      The EA-ADC(2) reduced density matrix is added to the MP2 reduced density matrix to form the state-specific density matrix.
@@
      The state-specific density matrix is diagonalized, and the resulting natural orbitals are sorted according to their occupation numbers.
@@
      Virtual natural orbitals with occupation numbers below the SS-FNO threshold are discarded.
@@
      EA-ADC(2) and EA-ADC(3) calculations are then performed in the truncated natural orbital basis and the corresponding truncated NAF basis.
@@
      A perturbative correction, defined as the difference between the canonical and truncated EA-ADC(2) results, is added to the EA-ADC(3) energy.
A schematic diagram of the SS-FNO-EA-ADC(3) algorithm is shown in FIG. 1.
All calculations in this work are performed using the density-fitting approximation. Core electrons were kept frozen for all calculations.
For complete basis set (CBS) extrapolation of electron affinities, the two-point cubic extrapolation scheme introduced by Helgaker and co-workers was employed. [25, 23]
@@
where $\alpha$ is a fitting parameter and $E_{CE}^{CBS}$ is the correlation energy at the CBS limit. Here, CE stands for correlation energy. The electron affinity results of aug-cc-pVXZ (X=D,T) were extrapolated to the CBS limit using Eq. (33).
@@ IV.1 Optimization of thresholds
To assess the performance of the SS-FNO scheme for the electron-attachment problem, we examined the deviation in EA-ADC(3) electron affinities as a function of the virtual space size in the canonical molecular orbital (MO) basis as well as in the natural orbital bases (FNO and SS-FNO) for the ozone molecule (see Fig. 2(a)). The experimental geometry of the ozone molecule [29] was used for the calculations. The aug-cc-pVQZ basis set and the aug-cc-pVQZ/C auxiliary basis set were used for the calculations. No NAF-based truncation was used in these calculations. The electron affinities (in eV) computed using different percentages of retained virtual orbitals are listed in the Supplementary Material (Table S1).
@@
Truncation based on occupation number thresholds is more robust than truncation based on a fixed percentage of retained virtual orbitals in natural-orbital-based wave-function methods. [30] In occupation-number-based truncation schemes, virtual natural orbitals with occupation numbers above a chosen threshold are retained in the calculations. Fig. 2(b) shows the convergence of the error in EA values at different natural-orbital truncation thresholds relative to the canonical EA-ADC(3) results. It can be seen that the SS-FNO-based scheme converges more rapidly than the corresponding FNO-based one. The EA values in the SS-FNO-based scheme converge a truncation threshold $10^{-5}$, while the FNO-based scheme shows a large error of -0.641 eV at the same threshold. This demonstrates that the SS-FNO framework is a more suitable approach for reduced-cost EA-ADC(3) calculations. The inclusion of the perturbative correction significantly enhances the accuracy, and the error converges at the threshold of $10^{-4}$. At this threshold, SS-FNO-EA-ADC(3) selects only $38.6\%$ of the total virtual space. Further reduction in computational cost can be achieved by reducing the auxiliary basis dimension using Natural Auxiliary Functions (NAFs). Fig. 3 illustrates the convergence of the error (in eV) in the EA value of O3 relative to the canonical reference, with decreasing NAF threshold and a fixed SS-FNO threshold of $10^{-4}$ in the EA-ADC(3) method. The EA values (in eV) at different NAF thresholds are listed in Table S3. The aug-cc-pVQZ basis set and the aug-cc-pVQZ/C auxiliary were used for the calculations. The convergence plot suggests that even at the NAF threshold of $10^{-1}$, the EA value converges with an error of 0.022 eV while retaining only about $70\%$ of the auxiliary basis functions. With the inclusion of perturbative corrections, the EA values show a negligible error (<0.01<0.01 eV) at the threshold of $10^{-1}$. However, for further benchmarking of different closed-shell systems, a conservative NAF threshold of $10^{-2}$ was chosen.
@@ IV.2 Benchmarking on EA24 test set
To further assess the performance of the SS-FNO-ADC(3) method, we calculated the electron affinities of the molecules included in the EA24 test set of Sherrill and co-workers. [58] This set consists of 24 organic acceptor molecules that possess bound valence-type electron-attached states, for which experimental values are available for most systems. The electron affinity values of the molecules in EA24 test set have been extensively studied and benchmarked for EA due to its practical implications in organic photovoltaics. [58, 20, 31, 13, 18, 66, 7, 51, 52, 41, 21] The molecules included in the EA24 test set are shown in Fig. 4, and their geometries were taken from Ref. 31. Some molecules possess more than one bound electron-attached state. However, we have considered only the lowest energy electron-attached state for the sake of simplicity.
@@
The vertical electron-attachment (EA) energies of these molecules were calculated at three truncation thresholds, $10^{-3}$, $10^{-4}$, and $10^{-5}$,using the aug-cc-pVDZ basis set together with the aug-cc-pVDZ/C auxiliary basis set. Fig. 5 shows the distribution of errors relative to the canonical EA-ADC(3) results. The plot shows a smooth convergence of the EA values as the truncation threshold is tightened. Furthermore, the inclusion of perturbative corrections leads to a noticeable reduction in both the spread and magnitude of the errors. The corresponding statistical parameters are summarized in Table 1using the following measures: MAD (maximum absolute deviation), ME (Mean Error), MAE (mean absolute error), STD (standard deviation), and RMSD (root mean squared deviation).The MAD value decreases from 0.378 eV at the $10^{-3}$ threshold to 0.042 eV at $10^{-4}$, and further to 0.003 eV at $10^{-5}$. The inclusion of the perturbative correction substantially improves the results, particularly at higher thresholds. For instance, at the $10^{-3}$ threshold, the MAD reduces from 0.378 eV to 0.170 eV upon applying the perturbative correction, while at $10^{-4}$, it decreases from 0.042 eV to 0.025 eV. At the tightest threshold ($10^{-5}$), the perturbative correction has a negligible effect, indicating convergence with respect to the SS-FNO truncation. The ME values suggest that the SS-FNO truncation systematically underestimates the EA, as reflected by the identical magnitudes of ME and MAE. All statistical parameters become negligible at the $10^{-5}$ threshold, confirming near-complete convergence of the SS-FNO approach.
Table 1: Statistical analysis of errors (in eV) of SS-FNO-EA-ADC(3) with respect to the canonical results at different FNO thresholds.
@@
Figure 6: Percentage of retained SSFNOs and absolute errors in EA values (eV) relative to canonical results, as a function of system size, for EA24 molecules using the aug-cc-pVDZ basis and aug-cc-pVDZ/C auxiliary basis
Fig. 6 shows the percentage of retained SSFNO (blue) and the absolute error in the EA energy (pink) as a function of system size. The percentage of retained SS-FNOs exhibits a variation ranging from approximately 65% to 85% across different systems. The truncation introduces a nearly uniform error across the test set, with the maximum absolute deviation remaining below 0.025 eV (including the perturbative correction). It demonstrates that the SS-FNO method can select an appropriate active space largely independent of the system size. The degree of truncation becomes even larger when larger basis sets are employed. For example, the average truncation for the EA24 test set is approximately 52 $\%$ in the aug-cc-pVTZ basis set (See Table S7). We also performed a statistical analysis of the errors arising from NAF truncation. Fig. 7 shows the distribution of errors relative to the canonical results for the SS-FNO-EA-ADC(3) method evaluated at different NAF truncation thresholds ($10^{-1}$ and $10^{-2}$). The perturbative correction is included for this comparison.
@@
Table 3: The electron affinities (in eV) of the first electron attached states of EA24 test set molecules calculated using various EA-ADC methods in CBS limit (in the SS-FNO basis, including perturbative correction) and compared with CCSD(T) results. The EA-EOM-DLPNO-CCSD/CBS results are also shown for comparison. The experimental values are provided for reference. The SS-FNO threshold is kept at $10^{-4}$ and NAF threshold at $0^{-2}. 11$footnotetext: NormalPNO with TCutPNOSingles == 1e-9 is used.
Molecule ADC(2)111in the canonical basis. $S$-FNO-ADC(3) $S$-FNO-sm-ADC[(2)+x(3)] EOM-DLPNO-CCSD22footnotemark: 2 CCSD(T) [58] Expt. [28]
@@
Table 3 summarizes the electron affinities (EAs) of the lowest electron-attached states for all molecules in the EA24 test set, computed using various EA-ADC methods and compared with the CCSD(T) reference values extrapolated to the complete basis set (CBS) limit. [58] The CBS extrapolation was performed using results obtained with the aug-cc-pVXZ (X = D, T) basis sets. For the second-order ADC calculations, canonical results are reported, whereas for the third-order ADC, only the corrected SS-FNO values are considered for comparison. The ADC(2) method exhibits the largest deviation from the CCSD(T) benchmarks, with a mean absolute deviation (MAD) of 0.973 eV. Inclusion of third-order corrections substantially improves the accuracy, as reflected in the reduced MAD of 0.56 eV for the SS-FNO-ADC(3) method, although this value remains larger than that observed for the corresponding EOM-DLPNO-CCSD method. The SS-FNO-sm-ADC[(2)+x(3)] approach demonstrates significantly improved performance, yielding a MAD of only 0.26 eV and showing close agreement with the EOM-DLPNO-CCSD results, which exhibit slightly smaller MAD and MAE but a larger standard deviation (STD). Although a direct comparison with experimental data is less meaningful due to the omission of vibrational contributions in the calculations, the SS-FNO-sm-ADC[(2)+x(3)] method at the CBS limit nonetheless exhibits marginally better agreement with experiment than the corresponding CCSD(T) results.